OpenStudy (anonymous):

what kind of math equation is 4b/2+3=11+b? Is it Algebra?

OpenStudy (anonymous):

Yes. Algebra

OpenStudy (anonymous):

how do I begin to solve this? what would be the steps?

OpenStudy (anonymous):

You have to try to get b on one side. If 3 is not in the denominator, then subtract 3 from both sides giving you 4b/2=8+b and 4/2=2, so you will now have 2b = 8+b, then subtract a b from both sides and you will have b = 8

OpenStudy (anonymous):

I think I understand. How would you know if 3 is not in the denominator? I tried to solve this by dividing 4 by 2 leaving 2b+3=11+b. I got stuck there.

OpenStudy (anonymous):

If 3 is on the bottom, then it is in the denominator, I can't tell if that's what you have since you don't have parentheses around it. You sure seem motivated at such a young age. Props to that!

OpenStudy (anonymous):

the equation is shown as this: ---4b/2+3=11+b.....it's a string so there isn't a "deominator" (I see what you're saying) How would I find out what "b" is after I do this: 2b+3=11+b

OpenStudy (anonymous):

2b+3 =11+b begin by rearranging the equation so that you have the unknowns on the left side of the "equal" signs and the known values on the right side of the "equal" sign. This is also known as making b the subject of the equation: thus: 2b +3 = 11 +b subtractract 3 from the left and the right side of the equation: 2b + 3 -3 = 11 +b -3 => 2b = 8 + b Now group b to the left hand side of the equation From the above equation we can acchieve this by subtracting b from both sides of the equation: Thus: 2b = 8 + b => 2b -b = 8 + b -b => b = 8 Therefor the answer is 8: A clean workout would like this: Beginning from your reduced equation: 2b + 3 = 11 + b => 2b = 8 +b ( after subtracting 3 from both sides) => b = 8 ( After subtracting b from both sides) The answer is b =8